OpenStudy (anonymous):

32 + (7 - 2) • 4 - 6 over 3

OpenStudy (anonymous):

Order of operations PEMDAS Parenthesis, multiplication, addition & subtraction from Left to Right (32 + 5 * 4 - 6)/3 can you take it from here

OpenStudy (anonymous):

22/3

OpenStudy (anonymous):

24.7 . << Is what I got ?

OpenStudy (anonymous):

32+20-6=46/3=46/3 or 15 1/3